OpenStudy (raden):

Find all pears (x,y,z) (real number) such that x^4 + y^4 + z^4 = 1

hero (hero):

Okay, this should work Let x = \(\sqrt[4]{\frac{1}{7}}\) y = \(\sqrt[4]{\frac{2}{7}}\) z = \(\sqrt[4]{\frac{4}{7}}\)

hero (hero):

Good luck finding the rest of them

hero (hero):

Me personally, I think it's infinite because you can always find three fractions that add to get 1

OpenStudy (anonymous):

If you want Integer solutions, there are only these : \[x = \pm1; y = 0; z = 0\] \[x = 0; y = \pm1; z = 0\] \[x = 0; y = 0; z = \pm1\] Observe that a number raised to the 4th power is always positive. Therefore only one variable can have the value 1, and the rest must be 0

hero (hero):

He never mentioned "integers" in the question, but okay.

hero (hero):

"all pairs of real numbers" doesn't imply "integers only"

OpenStudy (anonymous):

Yes, I know. I'm just giving additional info :)

hero (hero):

By the way, those numbers I posted can also be negative

hero (hero):

@RadEn, do you have anything to add or not. I'm about to hit the hay

OpenStudy (raden):

sorry, my conections very very slow, lost lost lost ohh :( Ok thank u.. yea, i got.. so, i should find the combin the sum of 0<=x^4+y^4<=1, and z will be take them

hero (hero):

The numbers can be less than zero, however, they will still be fractions

OpenStudy (anonymous):

to be precise : -1 <= x, y, z <= 1

OpenStudy (raden):

oh yea, negative number power even number must be positive... ok thank u

OpenStudy (anonymous):

one wonders how \((x,y,z)\) constitutes a pair

hero (hero):

family*